OpenStudy (anonymous):

Choose the value of the y determinant (Dy) in the following system. 2x – 4y = 24 2x – 3y = 15

OpenStudy (anonymous):

\[ \begin{bmatrix} 2&-4\\ 2&-3 \end{bmatrix} \begin{bmatrix} x\\y \end{bmatrix} = \begin{bmatrix} 24\\15 \end{bmatrix} \]

OpenStudy (anonymous):

To get the \(y\) determinant, first swap the \(y\) column with the product column\[ \begin{bmatrix} 2&24\\ 2&15 \end{bmatrix} \]

OpenStudy (anonymous):

Then find the determinant.

OpenStudy (anonymous):

\[ 2(15)-2(24) \]

OpenStudy (anonymous):

the choices are -132, -12, -18, and 78

OpenStudy (anonymous):

I'm telling you how to do it, not the answer.

OpenStudy (anonymous):

Learn, learn it

OpenStudy (anonymous):

-18 is the answer

OpenStudy (anonymous):

Did you learn how to do it?